Compute all the minors and cofactors of

1 2 3

2 0 1

2 3 4


1
Expert's answer
2021-05-26T03:27:19-0400

"A=\begin{bmatrix}
 1 & 2&3 \\
 2 & 0&1 \\
 2&3&4
\end{bmatrix}"


Let "M_{ij}" stands for Minors and "C_{ij}" is cofactors

where "i=i^{th}" row and "j=j^{th }" column of the matrix


"M_{11}=\begin{vmatrix}
 0 & 1 \\
 3 & 4
\end{vmatrix}=0-3=-3\ \ \ ;\ \ C_{11}=(-1)^{1+1}M_{11}=-3\\\ \\M_{12}=\begin{vmatrix}
 2 & 1 \\
 2 & 4
\end{vmatrix}=8-2=6\ \ ;\ \ C_{12}=(-1)^{1+2}M_{12}=-6\\\ \\M_{13}=\begin{vmatrix}
 2 & 0 \\
 2 & 3
\end{vmatrix}=6-0=6\ \ ;\ \ C_{13}= (-1)^{1+3}M_{13}=6\\\ \\M_{21}=\begin{vmatrix}
 2 & 3 \\
 3 & 4
\end{vmatrix}=8-9=-1\ \ ;\ \ C_{13}=(-1)^{2+1}M_{21}=1\\\ \\M_{22}=\begin{vmatrix}
 1 & 3 \\
 2 & 4
\end{vmatrix}=4-6=-2\ \ ;\ \ C_{22}=(-1)^{2+2}M_{22}=-2\\\ \\M_{23}=\begin{vmatrix}
 1 & 2 \\
 2 & 3
\end{vmatrix}=3-4=-1\ \ ;\ \ C_{23}=(-1)^{2+3}M_{23}=1\\\ \\M_{31}=\begin{vmatrix}
 2 & 3 \\
 0 & 1
\end{vmatrix}=2-0=2\ \ ;\ \ C_{31}=(-1)^{3+1}M_{31}=2
\\\ \\M_{32=}\begin{vmatrix}
 1 & 3 \\
 2 & 1
\end{vmatrix}=1-6=-5\ \ ;\ \ C_{32}=(-1)^{3+2}M_{32}=5\\\ \\M_{33}=\begin{vmatrix}
 1 & 2 \\
 2 & 0
\end{vmatrix}=0-4=-4\ \ ;\ \ C_{33}=(-1)^{3+3}M_{33}=-4"